# Trustworthy Semantic-Enabled 6G Communication: A Task-oriented and Privacy-preserving Perspective

### v1

@@ IV Experiments and Discussions
Two experimental evaluations (i.e. image classification and face recognition) are provided to assess both the task performance and the degree of privacy preservation within end-to-end ToSC systems, along with the attacker settings to improve the overall privacy-preserving process.
@@ IV-A Experimental Settings
1) Dataset and Attack Setups: CIFAR-10 dataset and CelebA dataset are adopted for image classification task and face recognition task, respectively. The former comprises $60,000$ color images, each measuring $32\times 32$ pixels, and categorized into $10$ distinct classes. The latter contains over 200,000 celebrity images with 40 attribute annotations per image, whose images cover a rich range of human pose variations and complex and diverse background information. To ensure comparability, all algorithmic networks under consideration are configured to produce outputs of identical dimensions. Furthermore, we incorporate a hypothetical scenario involving an adversarial attack network. This network is designed to execute model inversion via a black-box attack approach. It is posited to have continuous access to the network model on the target device, thereby enabling it to attempt image reconstruction. This scenario is pivotal in assessing the robustness of our framework against potential security breaches.
@@
2) Performance Metrics: For the image classification task, classification accuracy and PSNR of the attacker are employed. The former serves as an indicator of inference performance, with higher classification accuracy signifying more effective inference capabilities. The latter, the PSNR value, is utilized to gauge the level of privacy protection. In this context, a lower PSNR value in the reconstructed images indicates that its attacker steals the transmitted data, having a worse distortion of the reconstructed image, which also reflects the stronger privacy preservation, as it indicates that the reconstructed image has a reduced clarity or fidelity, which prevents unauthorised interpretation. For the face recognition task, the top-$1$ accuracy and the reconstructed image of the attacker are used. The former is a metric of recognition performance and is used to judge the task performance. The latter, i.e., the direct reconstruction effect of the image, is used to measure the level of privacy protection.
3) Approaches for Evaluation: To comprehensively demonstrate the efficacy of diverse privacy-preserving approaches, we have meticulously selected four state-of-the-art ToSC schemes, including:
@@
  • DeepJSCC-DP: DeepJSCC, originally designed for data-oriented communication systems, utilizes deep neural network-based encoders to map data directly to channel input symbols. To facilitate a comparative analysis of both performance and security aspects, we propose the integration of differential privacy mechanism [12], specifically through the injection of Laplacian noise into the transmission characteristics. The differential privacy mechanism allows for precise control over the level of privacy by adjusting the privacy budget, which is set at $0.05$, $0.1$, and $0.9$ for our experiments. A pivotal aspect to note is that a smaller privacy budget correlates with a higher volume of noise injected into the transmitted features. This increased noise level consequently leads to stronger privacy protection, as it more effectively obscures the original data features, thereby enhancing the security against potential data breaches or unauthorized data reconstruction efforts.
@@
  • DeepJSCC-Encryption: In DeepJSCC-Encryption, the encoder not only processes the data to extract features for the JSCC, but also integrates an encryption operation into its output [10]. At the encoder, encryption operations are performed on the encoded semantic features. This dual-functionality approach effectively combines feature extraction and encoding with a layer of cryptographic security. At the receiver, the process is reversed. The encoded and encrypted features are subjected to a decryption operation, a critical step for regaining the original data characteristics. Post-decryption, these features are then utilized for the intended classification and reconstruction tasks. This mechanism ensures that the data remains secure during transmission, only becoming accessible and usable upon successful decryption at the intended destination.
@@
  • IBAL: IBAL [11] represents a novel privacy-preserving scheme within the context of ToSC, leveraging the principles of adversarial learning. This method uniquely trains the encoder to effectively deceive the potential adversaries. It does so by optimizing the encoder to maximize the distortion in the data reconstruction process. Such a strategy is designed to thwart unauthorized attempts at data reconstruction, thereby enhancing the privacy and security of the transmitted information.
@@
  • DeepJSCC-LBVQ: DeepJSCC-LBVQ [13] represents a sophisticated ToSC scheme that incorporates digital modulation. Central to this approach is the implementation of a robust encoding framework, which is underpinned by a learned codebook. Its primary objective is to enhance communication robustness in response to channel variations. The essence of DeepJSCC-LBVQ lies in its ability to effectively balance the trade-off between informativeness and robustness. By employing a learned codebook, the scheme adapts to varying channel conditions, ensuring that the integrity and reliability of the transmitted data are maintained, even in challenging communication environments. This adaptability makes it a significant contender in scenarios where channel variability is a critical factor.
@@
The inference performance and the quality of image reconstruction are critically influenced by the dimensionality of the encoded representation. To facilitate fair and equitable comparisons across the different methods, we have standardized the dimensionality of the representations encoded by all methods under consideration. For the continuous representation methods such as DeepJSCC-DP, DeepJSCC-Encryption, and IBAL, we employ a full-resolution constellation modulation technique. This approach is instrumental in maintaining the integrity and resolution of the encoded data during the modulation process. In contrast, for DeepJSCC-LBVQ, which is a discrete method, we utilize the $M$-ary quadrature amplitude modulation (QAM) scheme. This choice is tailored to suit the discrete nature of the representations encoded by DeepJSCC-LBVQ, ensuring that the modulation process is compatible with the encoding method. Additionally, to further ensure the impartiality of our evaluation, we have standardized the settings across all adversary attack networks. This uniformity in settings is vital for ensuring that each method is subjected to non-discriminatory attacks, thereby providing a more accurate and fair assessment of private level. By adopting this approach, we aim to offer a comprehensive and unbiased comparison of each method’s ability to protect privacy under equivalent adversarial conditions.
@@ IV-B Information Leakage Under Adversarial Attack
Regarding the adversarial’s description, we assume ToSC system is under black-box model inversion attacks [14], where the adversaries reconstruct the received features as raw input using DNNs, and obtain users’ privacy. Specifically, the adversary network is a DNN designed by the adversary and deployed on the adversary’s device. And the transmitter’s coded features are illegally accessed by the adversary. The adversary then attempts to generate an approximate reconstruction of the user’s data based on the stolen transmission data. To improve the system’s capability to combat attacks, we make a weak assumption that attacker knows the codebook and can continuously access the trained encoder. As the intentions of attacker are not known, we consider a universal loss function for training the attacker neural network, which is to minimize the distortion of reconstructed data. For image transmission task, the loss function for training attacker can be expressed as
$\mathcal{L}_{Attack}$ $=\mathcal{L}_{MSE}+\mathcal{L}_{PER},$ (1)
where $\mathcal{L}_{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{x}_{i}-\tilde{\boldsymbol{x}}_{i}\right\|^{2}$ represents the MSE with $N$ denoting the number of samples, $\mathbf{x}_{i}$ representing the $i$-th sample, and $\tilde{\mathbf{x}}_{i}$ standing for the reconstructed data samples by the attacker; $\mathcal{L}_{PER}=\frac{1}{N}\sum_{i=1}^{N}\left\|VGN({\boldsymbol{x}_{i}})-VGN(\tilde{\boldsymbol{x}}_{i})\right\|^{2}$ represents the perception loss with $VGN(\cdot)$ being the first three layers of the pretrained visual geometry group’s (VGG) neural network [15].
@@ IV-C Experimental Results – Image Classification
Each scheme undergoes training at a specific signal-to-noise ratio (SNR), denoted as $\operatorname{SNR}_{\text{train}}$, which is set at $12$ dB. The testing phase involves varying SNR levels, specifically at $\operatorname{SNR}_{\text{test}}$ values of 00 dB, $4$ dB, $8$ dB, $12$ dB, $16$ dB, and $20$ dB. To maintain consistency and fairness, both discrete and continuous, the dimension of the encoded representation is uniformly set to $128$. This uniformity ensures that any observed differences in performance are attributable to the scheme’s inherent characteristics rather than discrepancies in encoded representation size. Specifically, for DeepJSCC-LBVQ method, which employs a discrete approach, we use a codebook of size $16$. This size is selected as it offers a balance between complexity and performance.
@@
(b) Attacker’s PSNR.
Figure 3: Classification Performance versus SNR on the CIFAR-10 dataset.
The analysis of Figs. 3(a) and 3(b) reveals the significant insights into the image classification performance of various privacy-preserving schemes under an additive white Gassian noise (AWGN) channel using the CIFAR-10 dataset. First, the classification accuracies of IBAL and DeepJSCC-LBVQ are significantly higher under regular channel conditions (i.e. SNR $\geq$ 4dB) compared to DeepJSCC-DP and DeepJSCC-Encryption. Moreover, these two schemes also exhibit superior privacy protection capabilities than DeepJSCC-DP method. Under low SNR regimes (i.e. SNR $<$ 4dB), DeepJSCC-DP and DeepJSCC-Encryption show better robustness in terms of classification accuracy. At this point, DeepJSCC-LBVQ classification accuracy decays, due to the wide deflection of discrete features caused by the codebook indexing of the channel transmission, but still maintains good privacy preserving ability. And IBAL shows the best privacy-preserving capability, as it focuses on trade-off between both privacy and task, with a slight bias at low SNR.
Notably, DeepJSCC-DP, with its increased Laplacian noise injection, offers improved privacy protection at the cost of task performance. For instance, DeepJSCC-DP with a privacy budget of $0.05$ achieves similar privacy protection levels as DeepJSCC-LBVQ, but its classification accuracy falls behind by approximately $3$ to $4$ when SNR is greater than $4$dB. DeepJSCC-Encryption presents the best privacy protection among the compared methods. However, this comes at the expense of task performance, failing to strike an optimal balance between privacy and utility. This contrast highlights the superiority of the intrinsic security incorporation schemes (IBAL and DeepJSCC-LBVQ) over feature perturbation methods in achieving a well-balanced effect in both task performance and privacy preservation. The advanced IBAL scheme particularly stands out for its capability to improve privacy protection without significantly compromising task inference performance, achieving an optimal privacy-utility trade-off compared to the baseline methods. A key factor contributing to DeepJSCC-LBVQ’s superior task performance is that its discrete representation, enhanced by the learned codebook, contains more informative messages, thereby leading to better performance. This underscores the effectiveness of discrete representation encoding in ToSC systems.
Figure 4: CelebA: from left to right, original, DeepJSCC-DP-0.9, DeepJSCC-DP-0.1, DeepJSCC-DP-0.05, DeepJSCC-Encryption, IBAL, DeepJSCC-LBVQ
@@ IV-D Experimental Results – Face Recognition
Next, we investigate the images of the Celeba dataset obtained after reconstructing the transmitted signals of these schemes using model inversion attacks. The results demonstrated in Fig. 4 can be seen that the attacker reconstructs the image worse with IBAL and DeepJSCC-LBVQ. IBAL even affects the attacker’s judgement on the gender reconstruction of men and women, which brings great protection. Combined with the above analysis of performance it can be seen that IBAL and DeepJSCC-LBVQ improve privacy protection while ensuring better task inference performance. And as Laplacian noise increases, DeepJSCC-DP scheme improve their privacy protection, i.e., the attacker reconstructs a blurrier image. In addition, the DeepJSCC-encryption scheme also provides good privacy protection because it completely disrupts the order of features in the image transmission, while the attacker only steals the image data and cannot carry out the complete reconstruction process later. In a comprehensive analysis, IABL and DeepJSCC-LBVQ schemes achieve a better privacy-utility trade-off compared to the baseline scheme.
@@ V-A Balancing Utility, Efficiency, and Privacy
In ToSC systems, achieving a harmonious balance between utility, efficiency, and privacy is a multifaceted challenge. Utility is paramount, as the primary goal is to transmit data that is semantically relevant to the task at hand. However, ensuring utility often requires processing and transmitting detailed information, which can inadvertently expose sensitive data, thereby conflicting with privacy goals. Efficiency is crucial in the high-volume, high-speed environment of 6G networks. Efficient data transmission not only conserves bandwidth but also reduces latency, enhancing the overall performance of the network. Privacy is perhaps the most challenging aspect to balance. While feature perturbation methods like encryption provide robust security, they do not address the unique challenges of real-time, semantically-rich communication. Advanced strategies, although more adaptable, come with their own set of risks, such as the inexplicability and uncertainty of black-box deep learning. To achieve this balance, a multi-layered approach [7] is often required. This involves combining feature perturbation privacy methods with intrinsic security incorporation techniques, continuously adapting to the changing dynamics of the network and data. Regular audits and updates to the privacy protocols are also essential to respond to new threats and technologies.

### v2

@@ IV-A Experimental Settings
1) Dataset and Attack Setups: CIFAR-10 dataset and CelebA dataset are adopted for image classification task and face recognition task, respectively. The former comprises $60,000$ color images and categorized into $10$ distinct classes. The latter contains over 200,000 celebrity images with 40 attribute annotations per image, whose images cover a rich range of human pose variations and diverse background information. All networks are configured to produce outputs of identical dimensions. Furthermore, we incorporate a hypothetical scenario involving an adversarial attack network. This network is designed to execute model inversion via a black-box attack approach. It is posited to have continuous access to the network model on the target device, thereby enabling it to attempt image reconstruction. This scenario is pivotal in assessing the robustness of our framework against potential security breaches.
@@
2) Performance Metrics: For image classification task, classification accuracy and MI leakage are employed. The former serves as an indicator of inference performance, with higher classification accuracy signifying more effective inference capabilities. The latter, is utilized to gauge the level of privacy protection. A lower MI leakage value in the reconstructed images indicates that its attacker steals the transmitted data, having less privacy leakage of the reconstructed image, which also reflects the stronger privacy preservation, as it indicates that the reconstructed image has a reduced sensitive data details, which prevents unauthorised interpretation. For face recognition task, the top-$1$ accuracy and the reconstructed image of the attacker are used. The former is a metric of recognition performance and is used to judge the task performance. The latter is used to measure the level of privacy protection.
3) Approaches for Evaluation: To make a comprehensive comparison, we select four state-of-the-art ToSC schemes, including:
@@
  • DeepJSCC-DP: DeepJSCC, originally designed for data-oriented communication systems, utilizes deep neural network-based encoders to map data directly to channel input symbols. Through the injection of Laplacian noise into the transmission characteristics [10], DP mechanism allows for precise control over the level of privacy by adjusting the privacy budget, which is set at $0.05$, $0.1$, and $0.9$ for our experiments. A pivotal aspect to note is that a smaller privacy budget correlates with a higher volume of noise injected into the transmitted features. This increased noise level consequently leads to stronger privacy protection, as it more effectively obscures the original data features, thereby enhancing the security against potential data breaches or unauthorized data reconstruction efforts.
@@
  • DeepJSCC-Encryption: In DeepJSCC-Encryption, the encoder not only processes the data to extract features for the JSCC, but also integrates an encryption operation into its output [8]. This dual-functionality approach effectively combines feature extraction and encoding with a layer of cryptographic security. At the receiver, the process is reversed. The encoded and encrypted features are subjected to a decryption operation, a critical step for regaining the original data characteristics. Post-decryption, these features are then utilized for the intended classification and reconstruction tasks. This mechanism ensures that the data remains secure during transmission, only becoming accessible and usable upon successful decryption at the intended destination.
@@
  • IBAL: IBAL [9] represents a novel scheme by leveraging the principles of adversarial learning. This method uniquely trains the encoder to effectively deceive the potential adversaries. It does so by optimizing the encoder to maximize the distortion in the data reconstruction process. Such a strategy is designed to thwart unauthorized attempts at data reconstruction, thereby enhancing the privacy and security of the transmitted information. Moreover, variational feature encoding (VFE) [11] belongs to the theory based on information bottleneck (IB), which is similar to IBAL. The scheme incorporates its IB theory and we comparatively test the performance of strategies with IB and without adversarial learning in terms of utility as well as privacy preservation.
@@
  • DeepJSCC-LBVQ: DeepJSCC-LBVQ [12] represents a sophisticated ToSC scheme that incorporates digital modulation. Its essence is the implementation of a robust encoder, which is underpinned by a learned codebook. Its primary objective is to enhance communication robustness in response to channel variations. The essence of DeepJSCC-LBVQ lies in its ability to effectively balance the trade-off between informativeness and robustness. By employing a learned codebook, the scheme adapts to varying channel conditions, ensuring that the integrity and reliability of the transmitted data are maintained, even in challenging communication environments. This adaptability makes it a significant contender in scenarios where channel variability is a critical factor. In addition, utility-informativeness-security-based ToSC (UIS-ToSC) [6] employs a strategy that combines adversarial learning as well as VQ and can be attributed to the LBVQ group. This scheme efficiently solves the utility-informativeness-security trade-off issue inherent in ToSC systems.
@@
The inference performance and the quality of image reconstruction are critically influenced by the dimensionality of the encoded representation. To facilitate fair comparisons, we have standardized the dimensionality of the representations encoded by all methods. For the continuous representation methods such as DeepJSCC-DP, DeepJSCC-Encryption, and IBAL, we employ a full-resolution constellation modulation technique. This approach is instrumental in maintaining the integrity and resolution of the encoded data during the modulation process. In contrast, for DeepJSCC-LBVQ, which is a discrete method, we utilize the $M$-ary quadrature amplitude modulation (QAM) scheme. This choice is tailored to suit the discrete nature of the representations encoded by DeepJSCC-LBVQ, ensuring that the modulation process is compatible with the encoding method. Additionally, to further ensure the impartiality of our evaluation, we have standardized the settings across all adversary attack networks. This uniformity is vital for ensuring that each method is subjected to non-discriminatory attacks, thereby providing a fair assessment of private level.
@@ IV-B Information Leakage Under Adversarial Attack
Regarding the adversarial’s description, we assume ToSC system is under black-box MIAs [13], where the adversaries reconstruct the received features as raw input using DNNs, and obtain users’ privacy. Specifically, the adversary network is a DNN designed by the adversary and deployed on the adversary’s device. And the transmitter’s coded features are illegally accessed by the adversary. The adversary then attempts to generate an approximate reconstruction of the user’s data based on the stolen transmission data. To improve the system’s capability to combat attacks, we make a weak assumption that attacker knows the codebook and can continuously access the trained encoder. As the intentions of attacker are not known, we consider a universal loss function for training the attacker neural network, which is to minimize the distortion of reconstructed data. For image transmission task, the loss function for training attacker can be expressed as the sum of the average MSE loss and the average perception loss across all sampling images.
@@ IV-C Experimental Results – Image Classification
Each scheme undergoes training at a specific signal-to-noise ratio (SNR), denoted as $\operatorname{SNR}_{\text{train}}$, which is set at $12$ dB. The testing phase involves varying SNR levels, specifically at $\operatorname{SNR}_{\text{test}}$ values of $4$ dB, $8$ dB, $12$ dB, $16$ dB, and $20$ dB. To maintain fairness, both discrete and continuous, the dimension of the encoded representation is uniformly set to $128$, which ensures that any observed differences in performance are attributable to the scheme’s inherent characteristics rather than discrepancies in encoded representation size. Specifically, for DeepJSCC-LBVQ method, which employs a discrete approach, we use a codebook of size $16$. This size is selected as it offers a balance between complexity and performance.
@@
(b) MI Leakage.
Figure 3: Task-accomplishment Performance (i.e., image classification) versus Privacy-preserving Performance (i.e., MI Leakage) on the CIFAR-10 dataset.
The analysis of Figs. 3(a) and 3(b) reveals the significant insights into the image classification performance of various privacy-preserving schemes under an additive white Gassian noise (AWGN) channel using the CIFAR-10 dataset. First, the classification accuracies of IB group (IBAL and VFE) and VQ group (DeepJSCC-LBVQ and UIS-ToSC) are significantly higher under regular channel conditions (i.e. SNR $\geq$ 8dB) compared to DeepJSCC-DP and DeepJSCC-Encryption. Moreover, two type of schemes (VQ scheme and IB-based IBAL) also exhibit superior privacy protection capabilities than DeepJSCC-DP method. However, although the classification accuracy of VFE is excellent, the privacy preservation is slightly worse than our proposed scheme (MI $>$ 1.55). Under low SNR regimes (i.e. SNR $<$ 8dB), DeepJSCC-DP and DeepJSCC-Encryption show better robustness in terms of classification accuracy. At this point, the VQ group (DeepJSCC-LBVQ and UIS-ToSC) classification accuracy decays, due to the wide deflection of discrete features caused by the codebook indexing of the channel transmission, but still maintains good privacy preserving ability. And IBAL shows the best privacy-preserving capability, as it focuses on trade-off between both privacy and task, with a slight bias at low SNR.
Notably, DeepJSCC-DP, with its increased Laplacian noise injection, offers improved privacy protection at the cost of task performance. For instance, DeepJSCC-DP with a privacy budget of $0.05$ achieves similar privacy protection levels as DeepJSCC-LBVQ, but its classification accuracy falls behind by approximately $3$ to $4$ when SNR is greater than $8$dB. DeepJSCC-Encryption presents the best privacy protection among the compared methods. However, this comes at the expense of task performance, failing to strike an optimal balance between privacy and utility. This contrast highlights the superiority of the intrinsic security incorporation schemes (IBAL and DeepJSCC-LBVQ) over feature perturbation methods in achieving a well-balanced effect in both task performance and privacy preservation. The advanced IBAL scheme particularly stands out for its capability to improve privacy protection without significantly compromising task inference performance, achieving an optimal privacy-utility trade-off compared to the baseline methods. A key factor contributing to DeepJSCC-LBVQ’s superior task performance is that its discrete representation, enhanced by the learned codebook, contains more informative messages, thereby leading to better performance. This underscores the effectiveness of discrete representation in ToSC systems.
[…]
@@ IV-D Experimental Results – Face Recognition
Next, we investigate the images of the CelebA dataset obtained after reconstructing the transmitted signals of these schemes using MIAs. As shown in Fig. 4, the attacker reconstructs the image worse with IBAL and DeepJSCC-LBVQ. IBAL even affects the attacker’s judgment on the gender reconstruction of men and women, which brings great protection. Combined with the above analysis of performance, it can be seen that IBAL and DeepJSCC-LBVQ improve privacy protection while ensuring better task inference performance. As Laplacian noise increases, DeepJSCC-DP scheme improve their privacy protection, i.e., the attacker reconstructs a blurrier image. Additionally, the DeepJSCC-encryption scheme also provides good privacy protection because it completely disrupts the order of features in the image transmission, while the attacker only steals the image data and cannot carry out the complete reconstruction process later. Thus, IABL and DeepJSCC-LBVQ schemes achieve a better privacy-utility trade-off compared to other schemes.
IV-E Scalability Issue
The scalability of algorithms in large-scale networks is influenced by the number of devices, affecting design and performance. For DeepJSCC+DP, while more devices improve data utility, they also increase computational and communication costs. Dynamic privacy budget allocation and adaptive noise strategies are essential to address these challenges. Distributed differential privacy (DP) and privacy amplification techniques help optimize performance, reduce heterogeneity, and minimize leakage risks. For DeepJSCC+Encryption, device count impacts complexity and overhead, requiring distributed architectures, hardware acceleration, and efficient key management. The IBAL algorithm faces scalability issues, with higher computational complexity and training demands, where adversarial training mitigates reconstruction risks. For DeepJSCC+LBVQ, the number of devices affects codebook size and complexity, necessitating efficient management. Moreover, the dynamic offloading framework plays a crucial role in large-scale privacy-preserving ToSC networks by effectively addressing dynamic traffic demands, optimizing resource utilization, and ensuring quality of service, thereby demonstrating significant application potential. For instance, the integration of traffic-aware network slicing and adaptive computation offloading strategies [14] facilitates precise traffic prediction and optimized resource allocation, thereby providing enhanced flexibility and adaptability in dynamic network environments [15].
IV-F Potential Applications
These methods can optimize the transmission of semantically rich, compressed data while ensuring privacy in AR applications, especially in high-throughput, low-latency immersive environments. However, real-world AR deployment must tackle large-scale data, device heterogeneity, and dynamic content challenges. Privacy is also crucial in remote medical diagnostics with sensitive patient data. Evaluating system effectiveness using public datasets is a start, with future work focusing on low-latency, secure medical communication. For autonomous vehicles, real-time, reliable data exchange is vital. The proposed approaches can prioritize important data while preserving privacy, but network congestion and scalability issues need further study. Adapting to strict performance criteria and integrating with vehicle protocols like V2X is essential for this application.
@@ V-A Balancing Utility, Efficiency, and Privacy
In ToSC systems, balancing utility, efficiency, and privacy presents a complex challenge. Utility denotes the system’s effectiveness and efficiency in executing downstream tasks, ensuring tasks are completed accurately and promptly, thereby enhancing overall performance. Informativeness pertains to the volume of data transmitted, impacting transmission speed and bandwidth use. Security focuses on the system’s ability to safeguard user privacy during data transmission and processing. While utility is paramount for transmitting semantically relevant data, its pursuit can inadvertently compromise privacy by exposing sensitive information. Efficiency is critical in 6G networks’ high-volume, high-speed context, where conserving bandwidth and reducing latency are key to performance. Privacy, however, is the most difficult to balance. Although encryption offers robust security, it doesn’t meet the demands of real-time, semantically-rich communication. More advanced strategies, despite their adaptability, introduce risks like the inexplicability of black-box deep learning. Achieving this balance typically requires a multi-layered approach [6], combining feature perturbation with intrinsic security techniques, and adapting to network and data dynamics. Regular audits and updates to privacy protocols are also vital to address emerging threats and technologies.